# Benchmarking Generalizable Bimanual Manipulation: RoboTwin Dual-Arm Collaboration Challenge at CVPR 2025 MEIS Workshop

### v1

@@ 2.3.2 Results Overview
3 Excellent Solutions (Two Tracks)
Figure 9: Simulation Track Champion.
Figure 10: Real-World Track Champion.
We highlight the winning solutions from both the Simulation and Real-World Tracks, AnchorDP3 [27] by the JD-TFS Team and SEM [17] by the TSAIL-HRL Team. Interestingly, both approaches incorporate explicit 3D modalities, in contrast to many recent VLA-based methods that rely solely on multi-view 2D visual inputs. This suggests that leveraging structured 3D representations can significantly enhance policy learning by improving sample efficiency and providing a more grounded understanding of physical space.
3.1 Simulation Track: AnchorDP3 by JD-TFS Team
@@
3.2 Real-World Track: SEM by TSAIL-HRL Team
@@
The TSAIL-HRL team’s SEM [17] framework secured the championship in the Real-World Track of Challenge, which reconceptualizes robot manipulation as a joint problem of explicit 3D spatial reasoning and embodiment-aware control. Rather than treating perception and action as loosely coupled subproblems, SEM tightly integrates multi-view vision, depth cues, and full kinematic state into a unified diffusion policy. At the architectural level, SEM comprises a spatial enhancer that lifts 2D image features into a coherent 3D embedding space, a robot state encoder that captures the full joint-graph structure of the dual-arm system, and a diffusion-based action decoder that conditions on fused semantic and geometric representations. The SEM pipeline is shown in Fig. 12.
@@
Empirically, SEM demonstrates that these design choices translate into both robustness and efficiency. The spatial enhancer and robot state encoder each contribute significant gains in ablation studies, and their combination yields state-of-the-art success rates across single- and multi-task benchmarks in simulation and on real hardware. Despite its theoretical rigor, SEM remains computationally tractable: lightweight, task-specific encoders minimize overhead, while the diffusion decoder provides a unified decision process. The result is a manipulation policy that generalizes gracefully to novel object arrangements, maintains high accuracy in the face of perceptual noise, and achieves superior task performance with fewer demonstrations and reduced training time.
@@ 6.2 Robot Learning in Manipulation
Recent advances in robotic manipulation have led to a wide range of policy architectures tailored for specific tasks and embodiments. Many task-specific approaches [9, 26, 6, 8, 5, 16, 22, 14, 15, 24, 23] achieve strong single-task performance but struggle to transfer across different embodiments or generalize to novel task configurations.
@@ 8 Conclusion
In this report, we introduced the winning solutions from both the Simulation and Real-World Tracks—AnchorDP3 and SEM—and provided an in-depth analysis of the techniques behind their success. In addition, we highlighted valuable observations and lessons shared by participating teams across all competition stages, offering a broad perspective on the current capabilities and challenges in building robust and generalizable dual-arm manipulation systems.

### v2

@@ 2.3.2 Results Overview
3 Excellent Solutions
We highlight the winning solutions, AnchorDP3 [29] by the JD-TFS Team and SEM [17] by the TSAIL-HRL Team. Interestingly, both approaches incorporate explicit 3D modalities, in contrast to many recent VLA-based methods that rely solely on multi-view 2D visual inputs. This suggests that leveraging structured 3D representations can significantly enhance policy learning by improving sample efficiency and providing a more grounded understanding of physical space.
3.1 AnchorDP3 by JD-TFS Team
@@
3.2 SEM by TSAIL-HRL Team
@@
The TSAIL-HRL team’s SEM [17] framework demonstrated remarkable performance in the simulation rounds, which reconceptualizes robot manipulation as a joint problem of explicit 3D spatial reasoning and embodiment-aware control. Rather than treating perception and action as loosely coupled subproblems, SEM tightly integrates multi-view vision, depth cues, and full kinematic state into a unified diffusion policy. At the architectural level, SEM comprises a spatial enhancer that lifts 2D image features into a coherent 3D embedding space, a robot state encoder that captures the full joint-graph structure of the dual-arm system, and a diffusion-based action decoder that conditions on fused semantic and geometric representations. The SEM pipeline is shown in Fig. 10.
@@
Empirically, SEM demonstrates that these design choices translate into both robustness and efficiency. The spatial enhancer and robot state encoder each contribute significant gains in ablation studies. Despite its theoretical rigor, SEM remains computationally tractable: lightweight, task-specific encoders minimize overhead, while the diffusion decoder provides a unified decision process. The result is a manipulation policy that generalizes gracefully to novel object arrangements, maintains high accuracy in the face of perceptual noise, and achieves superior task performance with fewer demonstrations and reduced training time.
@@ 6.2 Robot Learning in Manipulation
Recent advances in robotic manipulation have led to a wide range of policy architectures tailored for specific tasks and embodiments. Many task-specific approaches [9, 28, 6, 8, 5, 16, 24, 14, 15, 26, 25, 20, 19] achieve strong single-task performance but struggle to transfer across different embodiments or generalize to novel task configurations.
@@ 8 Conclusion
In this report, we introduced the winning solutions—AnchorDP3 and SEM—and provided an in-depth analysis of the techniques behind their success. In addition, we highlighted valuable observations and lessons shared by participating teams across all competition stages, offering a broad perspective on the current capabilities and challenges in building robust and generalizable dual-arm manipulation systems.